The net revenue of a company rose 3.6 %, while inflation rate was 3.0 %. Calculate the relative real change in revenue as percentage.

1
Expert's answer
2022-02-14T16:10:48-0500

Final value-initial value/initial value


=3.6-3.0=0.6

0.6/3.0=2×100=200%